# Exact Simulation of Longitudinal Data from Marginal Structural Models

### v2

@@ 5.2 Extensions and computational complexity
Another important extension involves including multiple time-varying confounders, observed or unobserved, such that $\bm{L}$ becomes a vector of $p$ variables. The number of calculations required for the distributions of ¯$\bm{\bar{}}L_{k-1}$, and $F(\tilde{Y}_{k+1}\,|\,do(\bar{A}_{k}),Y_{k}=0)$ among survivors at $k$ scales proportionally with $p$. The total number of computations becomes $\sum_{k=0}^{K-1}(2pk+3p+2)=K(K+2)p+2K$. Thus, the computational complexity grows linearly with $p$. In the example presented in Section 5.3, we simulate data with two time-varying confounders.
All of these calculations require evaluations of $h$-functions and their inverses. For commonly used copulas with analytic $h$-functions, these evaluations are computationally efficient and fast.
5.3 Illustrative example
We use an example to demonstrate the proposed algorithm simulating longitudinal time-to-event data, following a setup similar to Seaman and Keogh (2024). We consider baseline variables $B$ and $C$, two time-varying covariates $L_{k}$ and $W_{k}$, and treatment $A_{k}$ at follow-up $k$. The time-to-event is $T$, with $Y_{k+1}=I(T<k+1)$ indicating failure status observed at the end of the $k$th interval. The auxiliary continuous variable $\tilde{Y}_{k+1}$ is defined as previously. The data-generating models for the baseline and time-varying covariates, and treatment status are detailed in Appendix D. We assume a Cox marginal structural model given by:
$\lambda(\tilde{y}_{k+1}\,|\,do(\bar{a}_{k}),c,y_{k}=0)=\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right),$ (15)
where $\lambda_{0}=\exp(-2)$ is the baseline hazard. We set $\beta_{C}=-0.5$ and vary the values of $(\beta_{A},\beta_{CA})$ to construct scenarios with different strengths of treatment: a strong effect scenario $(-0.5,0.3)$ and a weak effect scenario $(0.2,0.1)$. This parameterization assumes a constant hazard rate between follow-up visits, influenced by the treatment status at the most recent visit. This implies that given survival at follow-up time $k$, the incremental survival time $\tilde{Y}_{k+1}$ follows an exponential distribution with the rate equal to the hazard above:
$\tilde{Y}_{k+1}\sim\operatorname{Exp}(\lambda_{0}\exp\left(\beta_{C}\,c+\beta_{A}\,a_{k}+\beta_{CA}\,c\,a_{k}\right)).$
If $\tilde{Y}_{k+1}<1$ then $Y_{k+1}=1$, and 00 otherwise. We set the total number of follow-ups $K$ to 6. We set the copulas describing the association between $L_{j}$ ($j=0,\ldots,k$) and $Y_{k+1}$ to be Student t with parameter $\rho_{L}$ and 5 degrees of freedom, although each pairwise dependence could, in principle, be distinct. Similarly, we model the copulas capturing the dependence between $Y_{k+1}$ and each $W_{j}$ ($j=0,\ldots,k$), given $L_{k}$, as Gaussian with correlation $\rho_{W}$, though again we note that different copulas could be specified for each time point. We vary the values of $\rho_{L}$ and $\rho_{W}$ to construct scenarios with different strengths of confounding (see Table 6 in Appendix D). We simulate data for sample size $n=\{1000,2000,5000\}$.
Remark 5.3.
Table 7 (in Appendix D) reports the runtime required to simulate data under the setup for the large causal effect with medium confounding scenario. The simulations were carried out on a laptop using a single score without parallelization. Generating data for $n$ up to 10,000 units takes only a few seconds, while even for a very large sample size ($n=50,\!000$) and 11 time steps ($K=10$), the runtime remains under one minute. This computational efficiency is crucial for tasks such as method evaluation and sample size calculations in clinical trial design, where simulations must be repeated across various parameter values.
We contrast two approaches for fitting the Cox marginal structural model specified in (15): stabilized inverse-probability-weighting and unweighted estimation. Details on weight calculation are provided in Appendix D. Table 4 summarizes the bias of estimates from 500 replications, compared to corresponding Monte Carlo standard errors as a measure of uncertainty. As the sample size increases, generally, the biases of estimates for $\beta_{C}$, $\beta_{A}$ and $\beta_{AC}$ obtained with inverse probability weights converge to zero, with their absolute values substantially smaller than the corresponding Monte Carlo standard errors. In contrast, the unweighted estimator exhibits considerable bias, particularly for the estimate of the treatment effect, $\beta_{A}$, which persists even at $n=5000$. The magnitude of this bias far exceeds the Monte Carlo standard errors. The unweighted estimates are most biased under the high level of confounding by $L$ and $W$, since this is not adjusted for.
Level of confounding Effect size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted High $1000$ $-0.04(2.27)$ $-1.05(1.54)$ $-0.14(2.31)$ $3.73(1.73)$ $0.27(3.09)$ $0.49(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.07(1.68)$ $-1.03(1.06)$ $-0.02(1.63)$ $3.81(1.21)$ $0.02(2.09)$ $0.34(1.51)$ $5000$ $0.02(1.25)$ $-0.96(0.68)$ $0.10(1.08)$ $3.89(0.76)$ $-0.09(1.58)$ $0.24(0.97)$ $1000$ $-0.03(2.35)$ $-0.93(1.59)$ $0.02(1.99)$ $3.64(1.46)$ $0.03(2.78)$ $0.44(1.94)$ $\beta_{A}=0.2$ $2000$ $-0.11(1.60)$ $-0.93(1.11)$ $0.02(1.49)$ $3.65(1.18)$ $0.17(2.02)$ $0.49(1.51)$ $5000$ $-0.03(1.14)$ $-0.95(0.71)$ $0.04(1.01)$ $3.65(0.71)$ $0.07(1.36)$ $0.51(0.89)$ Medium $1000$ $-0.07(2.04)$ $-0.43(1.51)$ $-0.13(2.19)$ $1.90(1.73)$ $0.23(2.91)$ $0.08(2.22)$ $\beta_{A}=-0.5$ $2000$ $-0.02(1.60)$ $-0.35(1.01)$ $0.00(1.65)$ $1.95(1.22)$ $0.09(2.27)$ $-0.01(1.55)$ $5000$ $-0.09(1.10)$ $-0.42(0.64)$ $-0.01(1.21)$ $1.97(0.80)$ $0.07(1.59)$ $-0.02(1.03)$ $1000$ $-0.12(2.06)$ $-0.36(1.43)$ $-0.04(2.02)$ $1.75(1.54)$ $0.14(2.70)$ $0.09(1.92)$ $\beta_{A}=0.2$ $2000$ $-0.10(1.62)$ $-0.42(1.08)$ $0.03(1.49)$ $1.68(1.11)$ $0.09(2.09)$ $0.17(1.44)$ $5000$ $-0.03(1.04)$ $-0.39(0.66)$ $0.03(0.95)$ $1.78(0.68)$ $0.03(1.32)$ $0.14(0.88)$ Low $1000$ $0.01(1.82)$ $0.46(1.29)$ $-0.23(2.36)$ $-1.95(1.83)$ $0.21(2.90)$ $-0.20(2.20)$ $\beta_{A}=-0.5$ $2000$ $-0.10(1.32)$ $0.44(0.89)$ $-0.04(1.69)$ $-1.83(1.25)$ $0.07(2.13)$ $-0.30(1.59)$ $5000$ $-0.02(1.04)$ $0.47(0.58)$ $0.01(1.07)$ $-1.77(0.84)$ $0.03(1.46)$ $-0.32(0.99)$ $1000$ $0.04(1.98)$ $0.53(1.30)$ $0.05(1.98)$ $-1.88(1.46)$ $-0.03(2.65)$ $-0.31(1.88)$ $\beta_{A}=0.2$ $2000$ $-0.02(1.45)$ $0.43(0.99)$ $0.04(1.43)$ $-1.85(1.10)$ $0.01(1.87)$ $-0.26(1.38)$ $5000$ $-0.01(0.92)$ $0.43(0.59)$ $0.07(0.95)$ $-1.89(0.68)$ $-0.06(1.25)$ $-0.25(0.84)$
Table 5 presents coverage rates of the 95% confidence intervals. Confidence intervals for the weighted estimated, computed using both sandwich variance estimation and bootstrap resampling, achieve coverage close to the nominal level, with slight undercoverage. In contrast, the sandwich-based confidence intervals for the unweighted estimator perform poorly, especially for the treatment effect estimate, $\beta_{A}$.
Level of confounding Effect Size $n$ $\beta_{C}$ $\beta_{A}$ $\beta_{AC}$ W-S W-B UW W-S W-B UW W-S W-B UW High 1000 93.8% 95.2% 88.4% 93.4% 92.8% 42.2% 93.8% 93.8% 93.4% $\beta_{A}=-0.5$ 2000 94.0% 94.0% 84.4% 95.2% 95.6% 12.4% 95.6% 95.6% 94.6% 5000 93.2% 92.8% 70.2% 94.4% 93.4% 0.2% 94.2% 93.6% 93.8% 1000 92.8% 93.8% 91.8% 96.2% 95.8% 31.6% 93.4% 94.0% 95.6% $\beta_{A}=0.2$ 2000 95.2% 95.0% 87.6% 94.6% 95.0% 9.8% 94.8% 94.6% 92.2% 5000 93.6% 93.6% 72.4% 93.8% 93.0% 0.2% 94.0% 94.0% 90.0% Medium 1000 93.8% 94.2% 93.4% 94.8% 93.6% 79.6% 93.8% 94.0% 93.6% $\beta_{A}=-0.5$ 2000 93.0% 93.8% 92.4% 95.2% 94.6% 64.8% 92.2% 93.4% 96.2% 5000 95.4% 94.8% 90.0% 93.2% 92.6% 30.2% 93.0% 93.8% 93.6% 1000 94.2% 94.8% 95.0% 95.0% 93.8% 77.2% 94.2% 93.2% 95.8% $\beta_{A}=0.2$ 2000 94.6% 93.8% 91.6% 93.6% 93.2% 64.0% 92.8% 92.2% 94.2% 5000 95.4% 94.6% 91.6% 95.4% 95.2% 26.8% 94.8% 94.4% 95.4% Low 1000 95.4% 95.4% 93.0% 95.0% 95.2% 84.0% 95.2% 95.2% 94.6% $\beta_{A}=-0.5$ 2000 95.6% 96.4% 93.0% 95.8% 96.0% 71.6% 95.0% 95.6% 94.6% 5000 93.2% 93.2% 87.2% 94.2% 94.8% 42.6% 94.0% 94.4% 94.4% 1000 94.6% 95.6% 94.8% 96.4% 96.6% 77.4% 92.8% 94.8% 96.0% $\beta_{A}=0.2$ 2000 94.8% 95.6% 92.8% 94.4% 94.6% 60.6% 95.0% 95.2% 94.2% 5000 94.8% 94.2% 89.4% 95.0% 95.4% 20.0% 95.0% 94.2% 93.8%

### v3

@@ 5.1 Simulation algorithm
If additional time-varying confounders—either observed or unobserved—are incorporated, such that $\bm{L}$ becomes a $p$-dimensional vector, the number of iterations needed of (14), (15) and (16) scales proportionally with $p$. As a result, the total number of computations becomes $\sum_{k=0}^{K-1}(3pk+2p+3)=K(3K+1)p/2+3K$. Thus, computational complexity grows linearly with the number of confounders $p$. In the example presented in Appendix E, we replicate the simulation study by Seaman and Keogh (2024), simulating data featuring two time-varying confounders.
However, it should be noted that if an individual experiences failure at any time step, the simulation terminates early. Hence, these stated complexities represent the maximum possible computational cost per individual. All of these calculations require evaluations of $h$-functions and their inverses and Copula distribution functions. For commonly used copulas, such evaluations are computationally efficient.
5.2 Illustration: Concomitant Treatment
We illustrate how the proposed simulation framework can support sample size calculation under complex, real-world scenarios. To do so, we construct a synthetic study inspired by the SOUL trial (Semaglutide cardiOvascular oUtcomes triaL) (McGuire et al., 2023). The SOUL trial is a randomized controlled study assessing the effects of oral semaglutide, a glucagon-like peptide-1 receptor agonist (GLP-RA), on cardiovascular risk in individuals with type 2 diabetes, and cardiovascular and/or kidney diseases. The primary endpoint is time to first occurrence of major adverse cardiovascular events (MACE), analyzed under the intention-to-treat principle.
A design challenge in SOUL arises from the permitted use of concomitant medications such as SGLT2 inhibitors (SGLT2-i), and there is demonstrably more uptake among participants in the control arm. The uneven ‘drop-in’ treatment between the semaglutide and placebo arms may dilute the observed intention-to-treat effect, because both groups end up receiving some active treatment, blurring the distinction between them, and potentially undermining statistical power if unaccounted for during sample size calculation. Our goal is to show how the proposed simulation approach can capture such intercurrent events while preserving a predefined causal estimand.
We simulate data over seven half-yearly follow-up periods. Let $B\in\{0,1\}$ indicate baseline cardiovascular disease, $L_{k}$ denote HbA1c at time $k$, $A\in\{0,1\}$ the randomized treatment (semaglutide), $S_{k}\in\{0,1\}$ SGLT2-i use at time $k$ and $Y_{k}$ the binary survival status at time $k$. We assume perfect adherence to $A$ and model event times using a Cox marginal structural model:
$\lambda(\widetilde{y}_{k+1}\,|\,do(A,\bar{S}_{k}),Y_{k}=0)=\lambda_{0}\exp\left(\beta_{A}\,A+\beta_{S}\,S_{k}+\beta_{AS}\,AS_{k}\right),$ (17)
where $\lambda_{0}$ is a baseline hazard assumed to be constant between follow-up visits and the survival indicator is calculated as $Y_{k+1}=I(\widetilde{Y}_{k+1}<1)$. Baseline characteristics and model parameters are loosely based on the design assumptions of the SOUL trial (McGuire et al., 2023), including an annual MACE rate of 3.5% and a hazard ratio of 0.83 for semaglutide (i.e. $\beta_{A}=\log(0.83)$). We construct 10 simulation scenarios by varying both the prevalence (0%, 5%, 10% and 15% annual drop-in) in the placebo group and efficacy (hazard ratios of 0.95,0.9,0.85) of SGLT2-i use. These hazard ratios reflect estimates reported for SGLT2 inhibitors in McGuire et al. (2021). Full specifications of the data-generating model and scenario parameters are provided in Appendix D. For each scenario and trial sizes of $n=\{5,\!000,10,\!000,20,\!000\}$, we simulate 400 datasets and fit a Cox MSM using inverse probability of treatment weights with stabilized weights. Naïve unweighted estimators are also computed for comparison. For reference, the total randomized population in SOUL is 9,650.
Figure 3: Boxplots of the bias for $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by inverse probability weighting (IPTW) across different drop-in scenarios.
Figure 3 and Figure 4 display boxplots of the estimation biases for the coefficients $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by inverse-probability-weighted and unweighted Cox regression, respectively. Across all scenarios, the inverse-probability-weighted estimates exhibit biases that are centered at zero. As the sample size increases, the distribution of these biases becomes narrower and more tightly concentrated around zero, indicating consistency. In contrast, unweighted estimates show substantial and persistent bias across all scenarios, with no evidence of convergence toward zero as the sample size increases. Table 4 presents the empirical coverage: the inverse-probability-weighted method achieves coverage close to the nominal 95% level, whereas the unweighted approach yields empirical coverage well below the nominal level, reflecting the impact of the time-varying confounding introduced to the data.
Figure 4: Boxplots of the bias for $\beta_{A}$, $\beta_{S}$ and $\beta_{AS}$ by unweighted Cox regression across different drop-in scenarios.
Effect of SGLT2-i Prevalence in placebo $n$ $\beta_{A}$ $\beta_{S}$ $\beta_{AS}$ IPTW Unweighted IPTW Unweighted IPTW Unweighted Strong effect $5000$ $96.5\%$ $65.0\%$ $94.0\%$ $0.0\%$ $93.0\%$ $74.0\%$ 5% drop-in $10000$ $94.8\%$ $47.3\%$ $95.0\%$ $0.0\%$ $94.3\%$ $49.0\%$ $20000$ $95.0\%$ $18.5\%$ $95.0\%$ $0.0\%$ $95.0\%$ $25.8\%$ $5000$ $94.0\%$ $28.0\%$ $95.0\%$ $0.0\%$ $92.3\%$ $64.5\%$ 10% drop-in $10000$ $95.8\%$ $4.5\%$ $96.5\%$ $0.0\%$ $95.3\%$ $42.5\%$ $20000$ $96.5\%$ $0.0\%$ $95.8\%$ $0.0\%$ $96.3\%$ $16.3\%$ $5000$ $92.8\%$ $4.8\%$ $94.0\%$ $0.0\%$ $94.0\%$ $60.8\%$ 15% drop-in $10000$ $94.5\%$ $0.0\%$ $96.0\%$ $0.0\%$ $94.3\%$ $34.0\%$ $20000$ $93.0\%$ $0.0\%$ $95.0\%$ $0.0\%$ $94.0\%$ $7.5\%$ Moderate effect $5000$ $95.3\%$ $68.0\%$ $96.5\%$ $0.0\%$ $95.5\%$ $69.3\%$ 5% drop-in $10000$ $94.5\%$ $43.3\%$ $93.8\%$ $0.0\%$ $94.3\%$ $50.8\%$ $20000$ $95.3\%$ $17.5\%$ $96.0\%$ $0.0\%$ $93.8\%$ $25.0\%$ $5000$ $96.0\%$ $25.5\%$ $97.3\%$ $0.0\%$ $95.5\%$ $63.8\%$ 10% drop-in $10000$ $95.0\%$ $4.5\%$ $96.3\%$ $0.0\%$ $93.5\%$ $40.8\%$ $20000$ $95.8\%$ $0.3\%$ $96.3\%$ $0.0\%$ $95.3\%$ $14.8\%$ $5000$ $94.3\%$ $4.8\%$ $95.5\%$ $0.0\%$ $95.8\%$ $58.0\%$ 15% drop-in $10000$ $95.3\%$ $0.0\%$ $94.5\%$ $0.0\%$ $94.0\%$ $37.0\%$ $20000$ $92.0\%$ $0.0\%$ $95.3\%$ $0.0\%$ $94.0\%$ $10.0\%$ Weak effect $5000$ $96.3\%$ $65.5\%$ $95.3\%$ $0.0\%$ $93.5\%$ $63.5\%$ 5% drop-in $10000$ $96.0\%$ $42.8\%$ $97.0\%$ $0.0\%$ $95.3\%$ $45.3\%$ $20000$ $95.0\%$ $13.3\%$ $95.0\%$ $0.0\%$ $95.0\%$ $18.5\%$ $5000$ $96.0\%$ $26.8\%$ $96.0\%$ $0.0\%$ $94.5\%$ $59.8\%$ 10% drop-in $10000$ $96.3\%$ $5.0\%$ $96.5\%$ $0.0\%$ $95.0\%$ $37.0\%$ $20000$ $95.5\%$ $0.0\%$ $96.8\%$ $0.0\%$ $96.3\%$ $11.5\%$ $5000$ $93.3\%$ $5.3\%$ $95.8\%$ $0.0\%$ $93.8\%$ $57.5\%$ 15% drop-in $10000$ $93.3\%$ $0.0\%$ $93.5\%$ $0.0\%$ $92.5\%$ $34.8\%$ $20000$ $95.5\%$ $0.0\%$ $95.5\%$ $0.0\%$ $95.0\%$ $7.3\%$ $5000$ $96.0\%$ $96.0\%$- -- -- -- No drop-in $10000$ $94.5\%$ $94.5\%$- -- -- -- $20000$ $93.5\%$ $93.5\%$- -- -- --
Table 4: Empirical coverage rates of 95% confidence intervals of inverse-probability-weighted (IPTW) and unweighted Cox regression methods.
These results confirm that the simulated data adhere to the marginal structural model specified in (17), and that time-varying confounding is induced. We now demonstrate how the proposed simulation framework can inform trial design, specifically for sample size determination under realistic scenarios involving intercurrent events.
The SOUL trial was designed to evaluate the effect of semaglutide on CV outcomes under the intention-to-treat principle, contrasting the outcomes of individuals randomized to semaglutide ($A=1$) versus placebo ($A=0$). Accordingly, we fit the following Cox proportional hazards model:
$\lambda(t\,|\,A)=\lambda_{0}\exp(\beta_{ITT}A),$
where $t$ is time to the first occurrence of MACE and $\exp(\beta_{ITT})$ represents the hazard ratio associated with being assigned semaglutide, irrespective of SGLT2-i use. Table 5 reports the estimated intention-to-treat effects and empirical power based on a one-sided type I error rate of $0.025$. In the hypothetical scenario with no SGLT2-i use in either arm, the estimated intention-to-treat effect closely approximates the true hazard ratio for semaglutide (0.83). Under this scenario, a trial with 10,000 participants achieves approximately 85% power. However, in all other scenarios, the intention-to-treat estimates are attenuated due to drop-in and consequently reduce power. To maintain at least 80% power, our simulations suggest that a sample size of 20,000 is sufficient but may be unnecessarily large and resource-intensive. Conversely, a trial with 5,000 participants is consistently underpowered, even in the absence of drop-in. A sample size of 10,000 appears to strike a reasonable balance: it achieves adequate power in scenarios with weak SGLT2-i effects, but falls short when the effect of SGLT2-i is strong or if the prevalence of drop-in in the placebo arm exceeds 5%. For the moderate scenarios, where SGLT2 inhibitors have a moderate effect, 10,000 participants is borderline sufficient, but if drop-in rates in the placebo arm exceed 5%, power may fall below acceptable thresholds. These findings suggest that 10,000 participants may be an appropriate design choice, but the sample size should be adjusted based on the plausibility of drop-in and effect size assumptions.
The 10 scenarios considered here are illustrative. The proposed simulation framework is highly flexible. For example, users can vary the marginal effect of semaglutide by adjusting $\beta_{A}$, alter the baseline incidence rate, or introduce more time-varying covariates. Because the algorithm relies on analytic function evaluations, it is computationally efficient: simulating 10,000 individuals over seven time steps requires approximately 40 seconds on a single core. This efficiency makes it practical and feasible to conduct large-scale simulation studies using parallel computation to support power analyses and inform robust trial design.
Effect of SGLT2-i Prevalence in placebo n Estimate Hazard Ratio Power Strong effect $5000$ $-0.164$ $0.849\%$ $46.5\%$ 5% drop-in $10000$ $-0.165$ $0.848\%$ $74.3\%$ $20000$ $-0.169$ $0.845\%$ $97.5\%$ $5000$ $-0.153$ $0.858\%$ $43.0\%$ 10% drop-in $10000$ $-0.148$ $0.862\%$ $65.5\%$ $20000$ $-0.151$ $0.860\%$ $93.0\%$ $5000$ $-0.143$ $0.867\%$ $37.5\%$ 15% drop-in $10000$ $-0.143$ $0.867\%$ $63.8\%$ $20000$ $-0.135$ $0.874\%$ $87.8\%$ Moderate effect $5000$ $-0.176$ $0.839\%$ $52.0\%$ 5% drop-in $10000$ $-0.173$ $0.841\%$ $81.0\%$ $20000$ $-0.179$ $0.836\%$ $98.8\%$ $5000$ $-0.160$ $0.852\%$ $45.0\%$ 10% drop-in $10000$ $-0.164$ $0.848\%$ $75.8\%$ $20000$ $-0.161$ $0.851\%$ $97.8\%$ $5000$ $-0.161$ $0.851\%$ $46.0\%$ 15% drop-in $10000$ $-0.157$ $0.855\%$ $76.3\%$ $20000$ $-0.154$ $0.857\%$ $94.5\%$ Weak effect $5000$ $-0.174$ $0.841\%$ $52.0\%$ 5% drop-in $10000$ $-0.181$ $0.834\%$ $84.5\%$ $20000$ $-0.184$ $0.832\%$ $98.3\%$ $5000$ $-0.176$ $0.838\%$ $53.3\%$ 10% drop-in $10000$ $-0.180$ $0.835\%$ $82.8\%$ $20000$ $-0.176$ $0.839\%$ $98.3\%$ $5000$ $-0.182$ $0.834\%$ $55.0\%$ 15% drop-in $10000$ $-0.176$ $0.838\%$ $82.3\%$ $20000$ $-0.173$ $0.841\%$ $98.3\%$ $5000$ $-0.187$ $0.830\%$ $56.8\%$ No drop-in $10000$ $-0.183$ $0.833\%$ $85.0\%$ $20000$ $-0.187$ $0.830\%$ $99.5\%$
Table 5: Under the intention-to-treat principal, estimates of $\beta_{ITT}$, hazard ratios ($\exp{\beta_{ITT}}$), and power using a one-sided type I error rate of $0.025$ for various scenarios and sample sizes ($n$).
5.3 Competing risks
An important consideration in research concerning time-to-event outcomes is that of competing risks; these are events whose occurrence prevents the observation of the outcome of interest—for example, non-cardiovascular death in studies of MACE. Classical literature defines various statistical estimands for competing risk settings, including the marginal, cause-specific and subdistribution hazards (see, for example, Pintilie, 2007; Geskus, 2016; Young et al., 2020). These estimands differ in interpretation and causal relevance.
Our proposed simulation framework offers the flexibility to efficiently simulate data from marginal structural models that target the controlled direct effect:
$F(\widetilde{Y}_{k+1}\,|\,do(\bar{A}_{k}=\bar{a}_{k},\bar{D}_{k}=\bar{0}_{k}),Z=z,Y_{k}=0)=f(\bar{a}_{k},z;\alpha),$ (18)
where $\{D_{k}\}_{k\in\mathcal{K}_{+}}$ denote the sequence of competing events and $do(\bar{D}_{k}=\bar{0}_{k})$ represents a hypothetical intervention eliminating these events. This aligns with the definition of marginal hazard in Young et al. (2020). Under this framework, competing risks are essentially treated as equivalent to censoring. To generate the competing event process, two approaches can be considered. The first specifies a fully conditional distribution: $F(\widetilde{D}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k},Z,Y_{k}=0)$. Under this approach, the outcome process $Y_{1:K}$ is simulated according to Algorithm 2, targeting the causal margin defined in (18). Once simulation of the failure process is complete—suppose the trajectory terminates at time step $K^{\prime}$ when $Y_{K^{\prime}+1}=1$—the competing event process is then simulated for each $k\in\{0,\ldots,K^{\prime}\}$. At each of these time steps, $\widetilde{D}_{k+1}$ is drawn from the specified conditional distribution $F(\widetilde{D}_{k+1}\,|\,\bar{L}_{k},\bar{A}_{k},Z,Y_{k}=0)$. If a competing event occurs before time $K^{\prime}$, the individual’s trajectory is truncated at the corresponding time step. Censoring events, such as loss to follow-up, can also be simulated using this approach.
Alternatively, a marginal structural model may be specified for the competing event hazard, mirroring the approach used for the primary outcome:
$F(\widetilde{D}_{k+1}\,|\,do(\bar{A}_{k}=\bar{a}_{k},\bar{Y}_{k}=\bar{0}_{k}),Z=z,D_{k}=0)=g(\bar{a}_{k},z;\beta).$ (19)
Additionally, we impose the following conditional independence:
[…]
which states that, under intervention, and conditional on baseline and time-varying covariates, $\widetilde{D}_{k+1}$ and $\widetilde{Y}_{k+1}$ are independent. Graphically, this corresponds to the absence of a direct edge between $Y_{k+1}$ and $D_{k+1}$ in the underlying causal graph—an edge that would not be identifiable from observed data in any case.
If both the outcome and the competing event are simulated using MSMs, the simulation proceeds as follows:
  1. 1.
    Simulate baseline covariates $Z$. Set $Y_{0}=D_{0}=0$.
  2. 2.
    For time point $k=0,\dots,K$:
    1. (a)
      If $Y_{k}=0$ or $D_{k}=0$, simulate $L_{k},A_{k}$; otherwise, stop;
    2. (b)
      If $Y_{k}=0$, simulate $Y_{k}$ following Algorithm 2; otherwise set $Y_{k}$ to null;
    3. (c)
      If $D_{k}=0$, simulate $D_{k}$ following Algorithm 2, substituting $Y$ with $D$, along with the corresponding distributions and copulas; otherwise, set $D_{k}$ to null.
  3. 3.
    Truncate this individual’s trajectory at the first occurrence of either the outcome of interest or the competing event.
Notably, in step (b), the decision to simulate $Y_{k+1}$ depends solely on $Y_{k}$ irrespective of $D_{k}$ and vice versa for step (c). This reflects the fact that we target a controlled direct effect, as specified in (18) and (19), where the other event is conceptually eliminated through intervention.
This approach is particularly helpful in settings like the estrogen therapy example in Young et al. (2020). By explicitly specifying the direction and magnitude of the direct effects of treatment on both prostate cancer mortality and other-cause mortality, one can simulate data to understand how these effects jointly determine the total effect, i.e. the cause-specific hazard. This framework can inform trial design: clinicians can input plausible direct effect assumptions and compute the sample size required to achieve adequate power for detecting a treatment effect on the cause-specific hazard.